OpenStudy (anonymous):

find sin theta if (12,-5) is a point on the terminal side of theta.

OpenStudy (anonymous):

only need the hypotenuse of the triangle. pythagoras should give you 13

OpenStudy (anonymous):

\[\sqrt{5^2+12^2}=\sqrt{25+144}=\sqrt{169}=13\] then take the appropriate ratios to find all the trig functions of theta

OpenStudy (anonymous):

-5/13 or 12/13?

Directrix (directrix):

Which one? Think about what the sine ratio means.

Directrix (directrix):

Hint: opposite side over hypotenuse

Directrix (directrix):